OpenStudy (anonymous):

Solve for x: 3(x + 1) = −2(x − 1) − 4

OpenStudy (konradzuse):

Lets do this again Distribute 3(x+1) = 3x+3 -2(x-1) -4 = -2x+2 -4

OpenStudy (konradzuse):

woops I'm sorry....

OpenStudy (konradzuse):

3x+3 = -2x-2 -3 to both sides 3x = -2x -5 add 2x to both sides 5x = -5 divide by 5 x = -1

OpenStudy (konradzuse):

it's -2x not 2x.

OpenStudy (anonymous):

Oh okay then x=-1